# Breaking the Frame: Visual Place Recognition by Overlap Prediction

### v2

@@ Abstract
Visual place recognition methods struggle with occlusions and partial visual overlaps. We propose a novel visual place recognition approach based on overlap prediction, called VOP, shifting from traditional reliance on global image similarities and local features to image overlap prediction. VOP proceeds co-visible image sections by obtaining patch-level embeddings using a Vision Transformer backbone and establishing patch-to-patch correspondences without requiring expensive feature detection and matching. Our approach uses a voting mechanism to assess overlap scores for potential database images. It provides a nuanced image retrieval metric in challenging scenarios. Experimental results show that VOP leads to more accurate relative pose estimation and localization results on the retrieved image pairs than state-of-the-art baselines on a number of large-scale, real-world indoor and outdoor benchmarks. The code is available at https://github.com/weitong8591/vop.git.
@@ 1 Introduction
Visual Place Recognition (VPR) is defined as the task of identifying the approximate location where a query image was taken, given a certain tolerance level [30, 55, 14, 2, 20, 60, 56, 19, 21, 26, 7, 23, 58]. VPR methods have been widely used in computer vision and robotics [12, 10] problems, such as autonomous, unmanned aerial [32], terrestrial, and underwater vehicles [43], as well as recent AR/VR devices.
Typically, VPR is approached as an image retrieval problem, where the image to be localized is compared against a large database of posed images and, optionally, a 3D reconstruction of the scene. Then, the most similar images retrieved from the database are used to estimate the precise location of the input query image, e.g., via local feature detection and matching. The complexity of VPR stems from a variety of factors such as seasonal variations, changes in viewpoint, the presence of dynamic objects, illumination differences, occlusions, weather conditions, and large-scale environments, as discussed in the studies by Hong et al.[22], Doan et al.[15], and Subramaniam et al[50].
@@
Recent approaches implement VPR as an estimation of image similarity and retrieval through learned image embeddings generated by a feature extraction architecture augmented with an aggregation or pooling mechanism, such as NetVLAD [2], CosPlace [6], and AnyLoc [25]. In such frameworks, images from the database are retrieved by optimizing their similarity between the embeddings to the query image, e.g., cosine similarity. One notable drawback of these methods is their sensitivity to partial overlaps resulting from occlusions. In practice, occlusions can significantly reduce image similarity scores despite the unobstructed portions of the scene, potentially offering valuable cues for localization. Second, such methods optimize the similarity of the embeddings on positives closer than the negative samples, which is hard to train as there is no ground truth (GT) prior to indicating the exact degree of similarity. Instead, we build GT patch matches as supervised by the 3D reconstruction. In addition, several reranking-based VPR methods [20, 62] retrieve the most similar images from the database on the shortlist built by global embeddings. Those improves the performance of retrieving correct images but needs more storage .
Instead of relying on global similarities or local features, we frame the problem as overlap prediction. This enables us to methodically understand which parts of a particular image are visible without having to perform feature detection and matching. Also, focusing on overlaps makes VOP more robust to occlusions and provides a better understanding of the query image in relation to the database. VOP encodes individual patches rather than the entire image, thereby facilitating the efficient establishment of patch-to-patch correspondences. These matches are subsequently utilized within a voting mechanism to calculate an overlap score for potential database images, providing a nuanced metric for retrieval amidst complex visual environments.
In Fig. 1, we show two retrieval examples using the proposed Visual Place Recognition by Overlap Prediction (VOP) and AnyLoc [25], a state-of-the-art retrieval solution. They process a query image alongside two database images: one sharing the scene with the query albeit with occlusion, and another from a different environment. AnyLoc erroneously attributes higher similarity to the unrelated image. In contrast, our proposed approach successfully identifies the correct image and the areas likely to overlap.
@@
    A novel approach focusing on patch overlap assessment, exhibiting enhanced robustness to challenges such as occlusions. The proposed method, VOP, learns patch embeddings suitable for visual localization and implements a single radius search across image patches in the database.
@@
    The introduction of a robust voting mechanism for image retrieval shows improvement over conventional similarity-based methods with nearest neighbors.
@@
    Breaking the frame of looking for close database images in standard VPR methods by considering geometric applications and using specific evaluation metrics.
@@ 2 Related Work
Global retrieval methods usually integrate or combine handcrafted or learning-based local descriptors, such as SIFT [31] through the BOW model [39, 48], or employ alternative algorithms [54]. There is a growing trend towards learning-based techniques that devise improved methods for feature aggregation. NetVLAD [2] introduces an end-to-end trainable image representation extractor, employing weak supervision through soft assignment to cluster centers. Generalized Mean Pooling (GeM) [40] proposes a learnable pooling layer that generalizes both max and average pooling, with a CNN serving as the backbone. Additionally, AP-GeM [42] enhances ranking accuracy by minimizing a relaxed global mean Average Precision (mAP) over a set of images. A CNN-based solution, CosPlace [6], utilizes the CNN backbone and GeM pooling but approaches the place retrieval issue as a classification problem. MixVPR [1] transitions feature maps into a compact representation space through cascade MLP blocks. AnyLoc [25] emerges as a versatile method suitable for various image retrieval contexts, exploring CNN-based or Vision Transformer (ViT)-based foundational models and aggregation techniques to construct a vocabulary for searching in a large database. More recently, DINOv2-SALAD [24] finetunes DINOv2 and proposes a new aggregation method based on optimal transport and the relations between features and clusters.
Image reranking. Visual place recognition techniques have increasingly embraced a two-phase approach that pairs global retrieval with subsequent reranking based on local features. MultiVLAD [3] is proposed to use the maximum similarity between the query VLAD descriptors and the query as the matching score. Similarly, Razavian. et al. [47] rely on the average L2 distance of each query sub-patch to the reference image. Patch-NetVLAD [20], for instance, innovates by integrating NetVLAD with region-level descriptors and multi-scale patch tokens for enhanced geometric verification. It effectively merges the strengths of local and global descriptors, offering high resilience to changes in condition and viewpoint. RRT [52] and TransVPR [57] further this development by incorporating image-level supervision and attention-based mechanisms, respectively, focusing on spatial relationships and feature relevance within images. CVNet and $R^{2}$ormer [62] introduce methodologies for replacing traditional geometric verification with dense feature correlation and transformer-based reranking, focusing on the precise alignment of features and relevance of image pairs. SuperGlobal [46] simplifies the reranking process by relying on global features and K-nearest-neighbors aggregation, showcasing a move towards more scalable and computationally efficient VPR solutions. These advancements highlight a trend toward utilizing machine learning techniques, including deep learning and transformers, to refine the accuracy and efficiency of image retrieval.
@@ 3.1 Problem Statement
Local features are usually image-dependent and redundant with variable lengths. Instead, we focus on patch-level representations and redefine the matching score as the overlap between the images. It is simple and efficient to achieve patch-level descriptors with less dependency. Let us split image $I$ into a set $\mathcal{P}$ of rectangular patches by a uniform grid consisting of $n$ rows and columns. Given sets of such rectangular patches $\mathcal{P}_{i}$ and $\mathcal{P}_{j}$ in the two images, there are multiple ways to define overlap. Here we define a patch-level overlap score as follows:
@@
Figure 2: Patch matching examples in an image pair (column) with varying patch sizes (16, 64, and 256). The 256 patches from DINOv2 [35] are average pooled to 16 and 64 patches. In each column, the numbers inside the patches indicate which ones are matched. The color overlay is calculated by applying PCA to the patch embeddings.
@@ 3.2 Learning Patch Matching via Patch Similarities
Figure 3: The learning pipeline of the proposed Visual Overal Prediction (VOP) method is shown, with a frozen ViT backbone [35], trainable encoder head, and contrastive loss on the similarities. The input to the training is image pairs that are broken down into rectangular patches to learn patch-to-patch overlap scores.
Figure 4: The proposed Visual Overlap Prediction (VOP) method at inference. Given an input query image and an image collection representing the reference map, the images are broken down into rectangular patches, the frozen backbone [35] extracts patch-level features, which are then fed into an encoder network, obtaining the final embeddings. The image retrieval is performed by running a radius search in the embedding space for each query patch to find potential database matches. The final overlap scores are determined by robust voting on the formed “query patch”-to-“database patch” correspondences. In practice, the map embeddings are obtained in an offline.
The pipeline for learning paytch-level embeddings is visualized in Fig. 3. Given a pair of images $I_{i}$ and $I_{j}$, the problem is to learn embeddings $e_{p}$ and $e_{q}$ as defined previously. First, we obtain patch features by inputting the images into a Visual Transformer architecture. We use a frozen DINOv2 [35] backbone. These features are then fed into an encoder head, comprising one linear projection layer plus a few fully connected layers with GELU activation and a dropout layer to encode the centers of the patches. Given the corresponding patch pair $(p,q)$ as a positive sample, we train this encoder to predict similar embeddings for $p$ and $q$. To do so, we employ contrastive learning on the similarities over the trained embeddings.
@@
For each query patch $p$, we prepared a set of candidate patches $\{q,q_{1},q_{2},...,q_{n}\}$ for training, where $\{q_{1},q_{2},...,q_{n}\}$ act as $n$ negative samples, depicting patches that do not have a common field-of-view with the query patch. In our training pipeline, half of the samples have no overlap (negative samples), e.g., they depict different scenes. The other half are from the same scene as the query. In the samples, there are 30-90% negative patches, which are the ones not matchable with the query patch. We consider the loss in a patch-level manner. We train our model by optimizing the contrastive loss on the similarities of the embeddings. The contrastive loss on patch-to-patch matching pairs is defined as:
@@
To provide means for quick prefiltering, we add a classification ([CLS]) token to the employed Transformer network that is trained to distill global image embeddings based on the image overlap scores, shown in the ablations. These tokens are used to efficiently select a subset of potentially overlapping images given the query.
Constructing Training Data. Now, we introduce the visual supervision used to build GT labels for training. The ground-truth overlapping patches of the training image pairs are achieved following LightGlue [29] employing pixel-wise depth and relative pose. As shown in Eq. 2, the patch overlap is measured by counting the co-visible 3D points in the reconstruction. The ground-truth matching local patches are found in two steps: First, a set of dense pixel coordinates $c_{i}$ are detected in the first image $I_{i}$, and transformed to 3D points $X_{i}$. The obtained 3D point cloud is filtered by non-zero depths and transformed to the reference camera, i.e., projecting the keypoints from camera $i$ to $j$ by the given transformation matrix, $kp_{i2j}$. Given camera intrinsics, we project the 3D points $X_{i}$ to the camera planes and check for visibility. We make sure that only cycle-consistent points are kept by using the projected point, calculating its 3D coordinates, and back-projecting to the original image for a visibility check again. The ground-truth matches are supposed to contain the points with a consistent depth [29]. This procedure provides the initial set of positive and negative pairs of matching image patches.
Next, given the visible pixel matches, we check which patches they belong to and define the matchable patches by having at least one visible pixel correspondence. We assume one-to-one matching relations, i.e., the patch with the most patch matches in the image pair will be chosen if there are multiple. Note that the visual supervision is built mutually from the image to the reference and vice versa.
@@ 3.3 Image Retrieval by Overlap Scores
Given a query image $I_{i}$ and a set of database images ${I_{j}}_{j=1,...,N}$, our goal is to efficiently determine the top-$k$ images for $I_{i}$ maximizing overlap. To do so, in the offline mapping phase of the method, we pre-generate the patch embeddings of all images in the database.
The retrieval pipeline is visualized in Fig. 4. In the online phase, we obtain the patch and, optionally, the global embeddings of the query image. Next, for each query image patch, we perform a radius search in the database of patch embeddings to determine all potentially overlapping candidates. Setting a radius threshold $\epsilon$ is easy as the employed cosine similarity falls between 0 and 1. We tune this threshold on the validation set and use a fixed one in all our experiments. Note that k-nearest-neighbors algorithm is insufficient in our case as we would like to find all potentially overlapping patches to determine image overlaps. Such a radius search can be efficiently performed by standard bounding structures, e.g., kd-tree, and other implementation tricks. Optionally, we also use the global descriptor to prefilter the images to a shortlist.
@@
Weighted Voting. To highlight the distinguishable query patches, we propose to weight the query patches from each image using a bag of visual words approach [11]. Specifically, we apply TF-IDF [41] weights to all the query patches, assigning lower weights to patches that frequently appear in queries but are rare in the matched database patches. As illustrated below, it is preferable to have more neighbor patches if fewer images share the same neighborhood with the query patch. For each query patch, the weight $t_{i}$ is calculated as follows:
@@
where $n_{id}$ is the number of neighbor patches in database images, defined as the cosine similarities of the patch descriptors within a specified radius, and $n_{d}$ is the total number of patches in each image, i.e., 256 in our case, with a patch size of (14, 14) on the resized image size of 224. Parameter $N$ represents the total number of database images, and $n_{i}$ denotes the number of database images that contain at least one neighbor of the query patch. Patches that frequently appear are deemed less important and thus are down-weighted. Weight $t_{i}$ for each query patch is incorporated into the overlap score computation as follows: $t_{i}*\hat{O}(I_{i},I_{j})$. Note that the map weights are precomputed during the mapping phase.

### v3

@@ Abstract
Visual place recognition methods struggle with occlusion and partial visual overlaps. We propose a novel visual place recognition approach based on overlap prediction, called VOP, shifting from traditional reliance on global image similarities and local features to image overlap prediction. VOP proceeds co-visible image sections by obtaining patch-level embeddings using a Vision Transformer backbone and establishing patch-to-patch correspondences without requiring expensive feature detection and matching. Our approach uses a voting mechanism to assess overlap scores for potential database images. It provides a nuanced image retrieval metric in challenging scenarios. Experimental results show that VOP leads to more accurate relative pose estimation and localization results on the retrieved image pairs than state-of-the-art baselines on a number of large-scale, real-world indoor and outdoor benchmarks. The code is available at https://github.com/weitong8591/vop.git.
@@ 1 Introduction
Visual Place Recognition (VPR) is defined as the task of identifying the approximate location where a query image was taken, given a certain tolerance level [35, 59, 15, 3, 23, 66, 61, 22, 24, 31, 8, 28, 63]. VPR methods have been widely used in computer vision and robotics [13, 11], such as autonomous, unmanned aerial [37], terrestrial, and underwater vehicles [47], as well as recent AR/VR devices.
Typically, VPR is approached as an image retrieval problem, where a query image is compared against a large database of posed images and, optionally, a 3D reconstruction of the scene. Then, the most similar images retrieved from the database are used to estimate the precise location of the input query image, e.g., via local feature detection and matching. The complexity of VPR often stems from various challenges including changes in viewpoint, the presence of dynamic objects, illumination differences, occlusion, weather conditions, seasonal variations, and large-scale environments, as discussed in the studies by Hong et al.[27], Doan et al.[16], and Subramaniam et al[54].
@@
Recent approaches implement VPR as an estimation of image similarity and retrieval through learned image embeddings. Such embeddings are generated by a feature extraction architecture combined with aggregation or pooling mechanisms, such as NetVLAD [3], CosPlace [7], and AnyLoc [30]. In such frameworks, images from the database are retrieved by optimizing the similarity between their embeddings and the query image, e.g., cosine similarity. One notable drawback of these methods is their sensitivity to partial overlaps caused by occlusion. In practice, occlusion can significantly reduce image similarity scores despite the unobstructed portions of the scene, potentially offering valuable cues for localization. The second drawback lies in training, as they optimize the similarity of the embeddings on positives closer than the negative samples. This process is inherently difficult because ground truth (GT) data does not provide explicit information about the exact degree of similarity. Instead, we build GT patch matches as supervised by 3D reconstruction. In addition, several reranking-based VPR methods [23, 68] retrieve the most similar images from a shortlist of database images built by global embeddings. Such methods improve the accuracy of image retrieval but require more storage.
Instead of relying on global similarities or local features, we frame the problem as overlap prediction. This enables us to methodically understand which parts of a particular image are visible without requiring feature detection and matching procedures. Also, focusing on overlaps makes VOP more robust to occlusion and provides a better understanding of the relations between the query and database images. VOP encodes individual patches rather than the entire image, thereby facilitating efficient patch-to-patch correspondences. These matches are subsequently incorporated into a voting mechanism to calculate an overlap score for potential database images, providing a nuanced metric for retrieval in complex visual environments.
In Fig. 1, we present two retrieval examples using the proposed Visual Overlap Prediction (VOP) method and AnyLoc [30], a state-of-the-art retrieval solution. They process a query image alongside two database images: one sharing the scene with the query but with occlusion, and the other from a different environment. AnyLoc incorrectly assigns a higher similarity to the unrelated image. In contrast, our proposed approach successfully identifies the correct image and the areas likely to overlap.
@@
    A novel approach shifting focus from image similarities to patch overlap assessment, showing enhanced robustness to challenges, e.g., occlusion. The proposed method, VOP, learns patch embeddings suitable for visual localization and implements a single radius search across image patches in the database.
@@
    The introduction of a robust voting mechanism for image retrieval shows improvements over conventional similarity-based methods with nearest neighbors.
@@
    Breaking the frame of looking for close database images in standard VPR methods by considering geometric applications and their evaluation metrics.
@@ 2 Related Work
Global retrieval methods usually integrate or combine handcrafted or learning-based local descriptors, such as SIFT [36] through the BOW model [43, 52], or employ alternative algorithms [58]. There is a growing trend towards learning-based techniques that devise improved methods for feature aggregation. NetVLAD [3] introduces an end-to-end trainable image representation extractor, employing weak supervision through soft assignment to cluster centers. Generalized Mean Pooling (GeM) [44] proposes a learnable pooling layer that generalizes both max and average pooling, with a CNN serving as the backbone. Moreover, AP-GeM [46] enhances ranking accuracy by minimizing a relaxed global mean Average Precision (mAP) over a set of images. A CNN-based solution, CosPlace [7], utilizes the CNN backbone and GeM pooling but approaches place retrieval as a classification problem. MixVPR [2] transitions feature maps into a compact representation space through cascade MLP blocks. AnyLoc [30] emerges as a versatile method suitable for various image retrieval contexts, exploring CNN-based or Vision Transformer (ViT)-based foundational models and aggregation techniques to construct a vocabulary for searching in a large database. More recently, DINOv2-SALAD [29] finetunes DINOv2 and proposes a new aggregation method based on optimal transport and the relations between features and clusters.
Image reranking. Visual place recognition techniques have increasingly embraced a two-phase approach that pairs global retrieval with subsequent reranking procedure based on local features. MultiVLAD [4] is proposed to use the maximum similarity between the query VLAD descriptors and the query as the matching score. Similarly, Razavian. et al. [51] rely on the average L2 distance of each query sub-patch to the reference image. Patch-NetVLAD [23], for instance, innovates by integrating NetVLAD with region-level descriptors and multi-scale patch tokens for enhanced geometric verification. It effectively merges the strengths of local and global descriptors, offering high resilience to changes in conditions and viewpoints. RRT [56] and TransVPR [62] further this development by incorporating image-level supervision and attention-based mechanisms, respectively, focusing on spatial relationships and feature relevance within images. CVNet and $R^{2}$ormer [68] introduce methodologies for replacing traditional geometric verification with dense feature correlation and transformer-based reranking, focusing on the precise alignment of features and relevance of image pairs. SuperGlobal [50] simplifies the reranking process by relying on global features and K-nearest-neighbors aggregation, showcasing a move towards more scalable and computationally efficient VPR solutions. These advancements highlight a trend toward utilizing deep learning techniques and transformers, to refine the accuracy and efficiency of image retrieval.
@@ 3.1 Problem Statement
Local features are usually image-dependent and redundant with variable lengths. Instead, we focus on patch-level representations and redefine the matching score as the overlap between the images. It is simple and efficient to achieve patch-level descriptors with less dependency. Let us split image $I$ into a set $\mathcal{P}$ of rectangular patches by a uniform grid consisting of $n$ rows and columns. Given sets of such rectangular patches $\mathcal{P}_{i}$ and $\mathcal{P}_{j}$ in the two images, there are multiple ways to define overlap, e.g., a patch-level overlap score as follows:
@@
Figure 2: A patch matching example with a patch number of $8^{2}$. The 256 patches from DINOv2 [40] are average pooled to 64. The numbers inside the patches indicate which ones are matched. The color overlay is calculated by PCA on patch embeddings.
@@ 3.2 Learning Patch Matching via Patch Similarities
Figure 3: The training pipeline of the proposed Visual Overlap Prediction (VOP) model, including a frozen DINOv2 backbone [40] breaking down input images into rectangular patches, a trainable encoder head, and contrastive loss using patch-to-patch overlap supervision.
Figure 4: The proposed VOP at inference. Given an input query image and a reference image collection, the frozen backbone [40] extracts patch-level features, which are then fed into our trained encoder to obtain the final embeddings. For each patch in the query image, the radius neighbor search is performed in the embedding space. The final overlap scores are determined by robust voting on the formed “query”-to-“database” patch neighbors. In practice, the map embeddings are offline pre-calculated and saved.
The pipeline for learning patch-level embeddings is visualized in Fig. 3. Given a pair of images $I_{i}$ and $I_{j}$, the goal is to learn embeddings $e_{p}$ and $e_{q}$ as defined previously. First, we obtain patch features by inputting the images into a frozen DINOv2 [40] backbone.These features are then fed into an encoder head, comprising a linear projection layer plus a few fully connected layers with GELU [26] activation and a dropout layer to encode the centers of the patches. Given the corresponding patch pair $(p,q)$ as a positive sample, we train the encoder to predict similar embeddings for $p$ and $q$. To do so, we employ contrastive learning on the similarities over the learned embeddings.
@@
For each query patch $p$, we prepared a set of candidate patches $\{q,q_{1},q_{2},...,q_{n}\}$ for training, where $\{q_{1},q_{2},...,q_{n}\}$ act as $n$ negative samples, depicting patches that do not have a common field-of-view with the query patch. In our training pipeline, half of the samples have no overlap (negative samples), i.e., they depict different scenes. The other half are from the same scene as the query. In the positive image samples, there are 30-90% negative patches, which are the ones not matchable with the query patch. We consider the loss in a patch-level manner. We train our model by optimizing the contrastive loss on the similarities of the embeddings of patch-to-patch matches:
@@
To provide means for quick prefiltering, we add a classification ([CLS]) token to the employed Transformer network that is trained to distill global image embeddings based on the image overlap scores, shown in the ablations of the Supp. Mat. These tokens are used to efficiently select a subset of potentially overlapping images given the query.
Supervision. Now, we introduce the visual supervision built for training. The ground-truth overlapping patches of the training image pairs are achieved following LightGlue [34] employing pixel-wise depth and relative pose. As shown in Eq. 2, the patch overlap is measured by counting the co-visible 3D points in the reconstruction. The ground-truth matching local patches are found in two steps: First, a set of dense pixel coordinates $c_{i}$ are detected in the first image $I_{i}$, and transformed to 3D points $X_{i}$. The obtained 3D point cloud is filtered by non-zero depths and transformed to the reference camera, i.e., projecting the keypoints from camera $i$ to $j$ by the given transformation matrix, $kp_{i2j}$. Given camera intrinsics, we project the 3D points $X_{i}$ to the camera planes and check for visibility. We make sure that only cycle-consistent points are kept by using the projected point, calculating its 3D coordinates, and back-projecting to the original image for a visibility check again. The ground-truth matches are supposed to contain the points with a consistent depth [34]. This procedure provides the initial set of positive and negative pairs of matching image patches.
Next, given the visible pixel matches, we check which patches they belong to and define the matchable patches by having at least one visible pixel correspondence. We assume one-to-one matching relations, i.e., the patch with the most patch matches in the image pair will be chosen if there are multiple. Note that the visual supervision is built mutually from the query image to the reference and vice versa.
@@ 3.3 Image Retrieval by Overlap Scores
Given a query image $I_{i}$ and a set of database images ${I_{j}}_{j=1,...,N}$, our goal is to efficiently determine the top-$k$ images for $I_{i}$ maximizing overlap. To do so, we pre-generate the patch embeddings of all images in the database during the offline mapping phase of the method.
The retrieval pipeline is visualized in Fig. 4. In the online phase, we obtain the patch and, optionally, the global embeddings of the query image. Next, for each query image patch, we perform a radius search in the database of patch embeddings to determine all potentially overlapping candidates. We use the rounded median similarity over 100 random samples as the radius threshold. Note that k-nearest-neighbors algorithm is insufficient in our case as we would like to find all potentially overlapping patches to determine image overlaps. Such a radius search can be efficiently performed by standard bounding structures, e.g., kd-tree, and other implementation tricks. For efficiency, we also use the global descriptor to prefilter the images to a shortlist.
@@
Weighted Voting. To highlight the distinguishable query patches, we propose to weight the query patches from each image using a bag of visual words approach [12]. Specifically, we adapt TF-IDF [45] weights to all the query patches, assigning lower weights to patches that frequently appear in queries but are rare in the matched database patches. As illustrated below, it is preferable to have more neighbor patches if fewer images share the same neighborhood with the query patch. For each query patch, the weight $t_{i}$ is calculated as follows:
@@
where $n_{id}$ is accumulated over all patch neighbors in a one-to-many manner, i.e., not forced to find only one neighbor in each database image. Patch neighbors are defined as the cosine similarities of the patch descriptors within a specified radius. Parameter $n_{d}$ is the total number of candidate patches in all DB images, i.e., 256 * $N$ in our case, with a patch size of $14^{2}$ on the resized image size of $224^{2}$. Parameter $N$ represents the number of DB images, and $n_{i}$ denotes the number of database images that contain at least one neighbor of the query patch. Patches that frequently appear are deemed less important and thus are down-weighted. Weight $t_{i}$ for each query patch is incorporated into the overlap score computation by $t_{i}*\hat{O}(I_{i},I_{j})$. Note that the map weights are precomputed during the mapping phase.